OpenStudy (anonymous):

Amy and Elizabeth both work in sales. Each week Amy earns $18/h plus a $40 bonus, and Elizabeth earns $12/h plus a $100 bonus. How many hours must Amy and Elizabeth each work in a week in order to earn the same amount of money? For that many hours, how much money would each earn? They would each have to work hours(?) , and they would earn (?) $ .

OpenStudy (texaschic101):

A = 18h + 40 E = 12h + 100 A = E (they make the same) now sub 18h + 40 = 12h + 100 --- subtract 40 from both sides 18h = 12h + 100 - 40 --- subtract 12h from both sides 18h - 12h = 100 - 40 --- combine terms 6h = 60 -- divide both sides by 6 h = 10 check... A = 18h + 40 E = 12h + 100 A = 18(10) + 40 E = 12(10) + 100 A = 180 + 40 E = 120 + 100 A = 220 E = 220 SO if they both worked 10 hrs, they would both make $ 220